• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 08:06:03 on March31,2005

    Entry number 142869

    Followups:

    keyword=Owl Shift Summary

    Owl Shift Summary
    ------------------
    
    Started shift with question about the NA Plane1 L-PMT TDCs 26 and the ADC for 
    this channel because they seemed to have few hits.  Called RC, and he said not 
    to worry about it for the night - will call the experts in the morning.  HV 
    seems to be ok for this channel.  TDC/ADC 26 in online gui plots corresponds to 
    the N1 scaler 27 according to Peter M.
    
    00:10 - Production Run 3085 ended.
    
    00:15 - Left Q1 tripped because of liquid level fluctuation.
    
    Called for controlled access to reset.  
    
    00:20 - ended Compton run 10517.
    
    00:50 - Q1 ramped back up.  Asked MCC for 45 uA of beam.
    
    Started taking data again at 00:53.
    Began Compton Run 10518 at 00:55.
    
    Production runs on tilted Carbon
    
    Run 3086 - 00:53-01:36  200k events
    
    Run 3087 - 01:36-02:17  200k events
    
    Run 3088 - 02:18-03:04  200k events
    
    Run 3089 - 03:05-03:49  200k events
    
    During run 3089, current changed from 45 uA to 40 uA.
    
    Run 3090 - 03:49-04:37  201k events
    
    MCC called and will try to run 45 uA again.
    
    Run 3091 - 04:38-05:37  171k events
    
    Need to reboot ROC5 because deadtime was high (99%)  - rebooted TS2.
    
    Compton Run 10518 - end of run failed.  Restarted coda, and started run 10519 
    at 05:14.
    
    Run 3092 - junk -> testing DAQ.
    
    After doing data replay, of the following runs:
    
    Run 3093 - 05:46-06:30  200k events
    
    Run 3094 - 06:30-07:16  200k events
    
    Run 3095 - 07:16-08:02  192k events
    
    we noticed that the Big Bite histograms are empty.  We still have a problem 
    with ROC5.  So this data is probably useless.  I'm sorry!
    
    We are now working to fix the problem.